OpenStudy (anonymous):

5square root x^4 times 7 square root x^2 5square root x^4 times 7 square root x^2 @Mathematics

OpenStudy (anonymous):

\[\sqrt[5]{x^4} \times \sqrt[7]{x^2}\]

OpenStudy (anonymous):

= x^(4/5) * x^(2/7) = x ^(28/35 + 10/35) = x^(38/35)